OpenStudy (anonymous):

Simply the equation Please help me :) 3 _____ 4+√5

OpenStudy (shadowfiend):

The best way to simplify these is to multiply the top and bottom by the denominator (that will get rid of the square root).

OpenStudy (shadowfiend):

So you would do: \[\frac{3(4 + \sqrt{5})}{(4 + \sqrt{5})(4 + \sqrt{5})}\]

OpenStudy (anonymous):

how do i do that

OpenStudy (shadowfiend):

... Do what? The above is what you do. You have to multiply it out.

OpenStudy (anonymous):

how do i do that?

OpenStudy (shadowfiend):

... By multiplying?

OpenStudy (shadowfiend):

For example, \(3(4 + \sqrt{5})\) = \(3\times 4 + 3\times \sqrt{5}\).

OpenStudy (shadowfiend):

You can leave \(3\times \sqrt{5}\) as just \(3\sqrt{5}\).

OpenStudy (anonymous):

and can you stop deleting my post :/

OpenStudy (shadowfiend):

No. You are already getting helped here. Please do not repost a question that you are already receiving help on. If you do repost a question that *hasn't been answered* and you aren't receiving help on, please delete your original question.

OpenStudy (anonymous):

okay i only did that b/c i don't understand what your talking about !

OpenStudy (shadowfiend):

Ok, that's fine. We'll try and fix that. In an equation like the one you posted, `simplifying' involves removing square roots from the denominator.

OpenStudy (shadowfiend):

To do that, we have to multiply the denominator by its conjugate (the same thing, but minus the square root instead of plus the square root), which will get rid of the square root. But, if we just do that in the denominator, we've changed the value of the entire equation. So, we multiply *both* the top *and* the bottom by the same thing, which is the same thing as multiplying by 1 -- i.e., it doesn't change the overall value.

OpenStudy (shadowfiend):

For example, if I have 3/5 and I multiply by 10/10, I have 30/50. But that's the same thing as 3/5, it's just written differently.